# Lotus: Diffusion-based Visual Foundation Model for High-quality Dense Prediction

### v4

@@ 6 Conclusion and Future Work
In this paper, we introduce Lotus, a diffusion-based visual foundation model for dense prediction. Through systematic analysis and tailored diffusion formulation, Lotus finds a way to better fit the rich visual prior from pre-trained diffusion models into dense prediction. Extensive experiments demonstrate that Lotus achieves SoTA performance on zero-shot depth and normal estimation with minimal training data, paving the way of various practical applications. Please see the supplementary materials for our discussion about Applications (Sec. F) and Future Work (Sec. G).
@@ A.1 Implementation Details
We implement Lotus based on Stable Diffusion V2 (Rombach et al., 2022), with text conditioning disabled. During training, we fix the time-step $t=1000$. To optimize the model, we utilize the standard Adam optimizer with the learning rate $3\times 10^{-5}$. All experiments are conducted on 8 NVIDIA A800 GPUs and the total batch size is 128. For our discriminative variant, we train for 4,000 steps, which takes $\sim$8.1 hours, while for the generative variant, we extend training to 10,000 steps, requiring $\sim$20.3 hours.
@@ Appendix B Details of Direct Adaption
As illustrated in Fig. 4 of the main paper, our Direct Adaption means directly adapting the standard diffusion formulation for dense prediction task with minimal modifications. Specifically, starting with the pre-trained Stable Diffusion model, image x and annotation y are encoded using the pre-trained VAE encoder. Noise is added to the encoded annotation to obtain the noisy annotation $\mathbf{z_{t}^{y}}$ at noise level $t\in[1,T]$. The encoded image $\mathbf{z^{x}}$ is then concatenated with the noisy annotation $\mathbf{z_{t}^{y}}$ to form the input of the denoiser U-Net model. To handle this concatenated input, the U-Net input layer is duplicated (from 4 channels to 8 channels) and its original weights are halved as initialization, which prevents activation inflation (Ke et al., 2024). Direct Adaptation is optimized using the standard multi-step formulation the standard diffusion objective, $\epsilon$-prediction, as described in Eq. 2 (first row) of the main paper. To analyze the original diffusion formulation more effectively, we avoid specialized techniques introduced in prior methods (Ke et al., 2024; Fu et al., 2024; Xu et al., 2024; Ye et al., 2024), such as annealed multi-resolution noise and test-time ensembling.
@@ Appendix D Performance of $v$-prediction
Appendix E Qualitative Comparisons
@@
Appendix F Applications of Lotus
@@
Appendix G Future Work

### v5

@@ 6 Conclusion and Future Work
In this paper, we introduce Lotus, a diffusion-based visual foundation model for dense prediction. Through systematic analysis and tailored diffusion formulation, Lotus finds a way to better fit the rich visual prior from pre-trained diffusion models into dense prediction. Extensive experiments demonstrate that Lotus achieves promising performance on zero-shot depth and normal estimation with minimal training data, paving the way of various practical applications. Please see the supplementary materials for our discussion about Applications (Sec. I) and Future Work (Sec. J).
@@ A.1 Implementation Details
We implement Lotus based on Stable Diffusion V2 (Rombach et al., 2022), with text conditioning disabled. Both the depth and normal maps are normalized to the range $[-1,1]$ to match the designed input value range of the VAE. During training, we fix the time-step $t=1000$. To optimize the model, we utilize the standard Adam optimizer with the learning rate $3\times 10^{-5}$. All experiments are conducted on 8 NVIDIA A800 GPUs and the total batch size is 128. For our discriminative variant, we train for 4,000 steps, which takes $\sim$8.1 hours, while for the generative variant, we extend training to 10,000 steps, requiring $\sim$20.3 hours.
@@ Appendix B Details of Direct Adaption
As illustrated in Fig. 4 of the main paper, our Direct Adaption means directly adapting the standard diffusion formulation for dense prediction task with minimal modifications. Specifically, starting with the pre-trained Stable Diffusion model, image x and annotation y are encoded using the pre-trained VAE encoder. Noise is added to the encoded annotation to obtain the noisy annotation $\mathbf{z_{t}^{y}}$ at noise level $t\in[1,T]$. The encoded image $\mathbf{z^{x}}$ is then concatenated with the noisy annotation $\mathbf{z_{t}^{y}}$ to form the input of the denoiser U-Net model. To handle this concatenated input, the U-Net input layer is duplicated (from 4 channels to 8 channels) and its original weights are halved as initialization, which prevents activation inflation (Ke et al., 2024). Direct Adaptation is optimized using the standard multi-step formulation the standard diffusion objective, $\epsilon$-prediction, as described in Eq. 2 (first row) of the main paper. To analyze the original diffusion formulation more effectively, we avoid specialized techniques introduced in prior methods (Ke et al., 2024; Fu et al., 2024; Xu et al., 2024; Ye et al., 2024), such as annealed multi-resolution noise (AMRN).
Table 4: Experiments based on Marigold $w/$ AMRN.
Index Method NYUv2 KITTI
[…]
1-1 $\epsilon$-pred. 6.746 95.021 11.827 87.065
1-2 $\epsilon$-pred. + single step 6.691 94.552 13.395 76.269
1-3 $\epsilon$-pred. + single step + detail preserver 6.547 94.772 12.815 77.829
2-1 $v$-pred. 6.358 95.188 10.796 89.726
2-2 $v$-pred. + single step 5.499 96.415 11.132 88.520
2-3 $v$-pred. + single step + detail preserver 5.422 96.517 10.761 89.826
3-1 $x_{0}$-pred. 6.262 95.501 10.769 89.643
3-2 $x_{0}$-pred. + single step 5.495 96.431 11.237 88.457
3-3 $x_{0}$-pred. + single step + detail preserver 5.418 96.542 10.651 89.887
Table 5: Experiments based on Marigold $w/o$ AMRN.
Index Method NYUv2 KITTI
[…]
1-1 $\epsilon$-pred. 13.110 85.083 17.655 75.581
1-2 $\epsilon$-pred. + single step 6.605 94.583 13.406 76.298
1-3 $\epsilon$-pred. + single step + detail preserver 6.582 94.768 12.823 77.983
2-1 $v$-pred. 10.634 89.448 14.328 84.026
2-2 $v$-pred. + single step 5.498 96.562 11.173 88.314
2-3 $v$-pred. + single step + detail preserver 5.459 96.657 10.814 89.081
3-1 $x_{0}$-pred. 8.058 92.834 12.177 86.301
3-2 $x_{0}$-pred. + single step 5.477 96.615 11.166 88.640
3-3 $x_{0}$-pred. + single step + detail preserver 5.396 96.717 10.575 89.804
The AMRN strategy aims to reduce the model’s variance, which has a similar effect to our design, $x_{0}$-pred., but through a different solution. This diminishes the impact of our method. Therefore, it is preferable to validate the effect of our designs $w/o$ AMRN. We validate this claim using the Marigold codebase, both $w/$ and $w/o$ AMRN, as shown in the Tab. 4 and Tab. 5, respectively. In Tab. 5, the performance of multi-step models follows the order: $\epsilon$-pred. $<$ $v$-pred. $<$ $x_{0}$-pred. However, in Tab. 4, the differences between three parameterization types are minimal, particularly the performance of $v$-pred. and $x_{0}$-pred. are nearly identical. This can be attributed to the influence of AMRN, which is specifically designed for multi-step diffusion models to reduce variance and enhance performance. As a result, $x_{0}$-pred. shows no significant difference in reducing variance compared to the other two parameterizations. In Tab. 5, when the number of time-steps is reduced to one, the performance of the model improves regardless of the parameterization type used. However, in Tab. 4, the effect of single-step is unstable. This unexpected phenomenon arises from the complex, multifaceted effects of AMRN when transitioning from multi-step to single-step: ① AMRN significantly improves the multi-step model, but its effect is lost when the number of time-steps is reduced to one. ② In the single-step model, convergence is easier with limited data, leading to a slight improvement in performance. However, this also leads to catastrophic forgetting, which reduces the model’s ability to handle detailed areas, especially on the KITTI dataset. In both Tab. 4 and Tab. 5, Detail Preserver further enhances the performance of single-step model, particularly on the KITTI dataset, which contains more complex and detailed areas, such as pedestrians and fences, compared to the NYUv2 dataset. In both Tab. 4 and Tab. 5, when using a single step ($t=T$), according to $\mathbf{v}_{t}=\sqrt{\bar{\alpha}_{T}}\bm{\epsilon}-\sqrt{1-\bar{\alpha}_{T}}\mathbf{z}$, since $\sqrt{\bar{\alpha}_{T}}\approx 0$ when $t=T$, $v$-pred. becomes equivalent to $x_{0}$-pred. This explains why the performances of $v$-pred. and $x_{0}$-pred. are nearly identical in single-step, with only minor differences. In conclusion, these experiments show that AMRN, which has a similar effect to our designs but is achieved through a different solution, diminishing the impact of our proposed designs. Therefore, it is preferable to validate the effect of our designs $w/o$ AMRN. The experiments on Marigold $w/o$ AMRN (Tab. 5) validate the effectiveness of our proposed designs, as stated in our main paper, where the best protocol is $x_{0}$-pred. + single step + detail preserver.
@@ Appendix D Performance of $v$-prediction
Appendix E Experiments on More Dense Prediction Tasks:
       Semantic Segmentation and Diffuse Reflectance
Figure 12: Experiments of Lotus on (a) semantic segmentation and (b) diffuse reflectance. The high-quality results indicate that our method, even without task-specific designs, can be effectively applied not only to geometric dense prediction tasks, but also to semantic dense prediction tasks.
Table 6: The quantitative results of semantic segmentation on Hypersim (Roberts et al., 2021) testing set. Mean values are reported from 10 independent runs.
Method mIoU $\uparrow$ mAcc $\uparrow$
Direct Adaption 14.1 61.3
Lotus-G 21.2 65.6
Table 7: The quantitative results of diffuse reflectance prediction on Hypersim (Roberts et al., 2021) testing set. Mean values are reported from 10 independent runs.
Method L1 $\downarrow$ L2 $\downarrow$
Direct Adaption 0.198 0.206
Lotus-G 0.109 0.135
To validate the generalization ability of our method on other dense prediction tasks, we further train it on semantic segmentation and diffuse reflectance prediction. Both tasks are trained using the training set of the Hypersim dataset (Roberts et al., 2021) and evaluated on their corresponding test sets. For semantic segmentation, we report the mean intersection over union (mIoU) and mean accuracy (mAcc). For diffuse reflectance prediction, we evaluate using the L1 and L2 distances to the ground truth. To enable fast evaluation, we randomly select 500 paired testing samples. In our experiments, we do not redesign any specific modules or loss functions for these tasks and maintain the original training protocol of Lotus unchanged. As shown in Tab. 7 and Tab. 7, we compare our method with the baseline, Direct Adaption (Fig. 4 in the main paper), to assess its effectiveness. The results show that our method outperforms the baseline across all metrics. Additionally, we provide qualitative visualizations for these two tasks in Fig. 12, demonstrating accurate and high-quality results. Both the quantitative and qualitative results indicate that our method, even without task-specific designs, can be effectively applied not only to geometric dense prediction tasks, as shown in the main paper, but also to semantic dense prediction tasks.
Appendix F Frequency Domain Analysis of the Detail Preserver
         Take Monocular Depth Estimation as An Example
We use fast Fourier transform (FFT) to compute the Discrete Fourier Transform (DFT) of the input images and depth map estimations with and without Detail Preserver. The entire 2D frequency domains are divided into 8 frequency groups exponentially using the base of 2, i.e., the first group covers the 2D frequency map in a circle with a radius of 2, the second group covers the annular region with radii from 2 to 4, the third group covers radii from 4 to 8, and so on. This exponential grouping allows us to analyze the frequency components across progressively larger ranges, capturing both low-frequency and high-frequency characteristics.
(a) Frequency domain energy distribution comparisons among input image, and depth estimations w/ and w/o Detail Preserver.
(b) Frequency energy ratio between input image and GT depth.
(c) Frequency energy ratio between input image and depth estimations w/ Detail Preserver.
(d) Frequency energy ratio between the input image and depth estimations w/o Detail Preserver.
Figure 13: Frequency Domain Analysis of the Detail Preserver We use Hypersim (Roberts et al., 2021) dataset to transfer the input image and depth estimation w/ and w/o Detail Preserver into 2D frequency domains, using FFT. 100 pairs of {input image, depth estimation w/ Detail Preserver, depth estimation w/o Detail Preserver} are randomly selected for this frequency domain analysis. Hypersim is a photorealistic synthetic dataset. Not only can Hypersim offer dense GT labels without None areas (which is important during FFT), its depth annotations are much fine-grained compared with real-world datasets like NYUv2 (Silberman et al., 2012) and KITTI Geiger et al. (2013).
In order to more clearly demonstrate the effect of our proposed Detail Preserver, we first analysis the experiments using Hypersim (Roberts et al., 2021) dataset to display the difference in frequency domain energy between the details from both geometry and texture (the input images); and the details from purely the geometry (the GT depth maps). As shown in Fig. 13(b), the frequency domain energy between the input images and the depth annotations are plotted. Clearly we can see that the input images has much higher frequency energy in high-frequency areas, i.e., group 4, 5, 6, and 7, indicating that the details in surface textures mainly contribute to high-frequency energy; while the details in geometries, which can be expressed by depth maps, are mainly concentrated into (relative) middle and low frequency areas, i.e., group 0, 1, 2, and 3.
As shown in Fig. 13(a) , collaborating with the Detail Preserver effectively drag the frequency domain energy of depth estimation to the input image, especially on middle and low frequency domains, i.e., the frequency group 0, 1, 2 and 3, highlighting the Detail Preserver’s effectiveness in enhancing the geometrical details that should be reflected into depth predictions, like the fences around roads and houses (Fig. 8 of our main paper).While for high-frequency components, i.e., the frequency group 4, 5, 6, and 7, which may be primarily caused by the highly detailed textures, like the signs on the road and patterns on house surfaces, the energy in these areas between depth estimations with and without Detail Preserver is quite similar, indicating that the Detail Preserver does not copy this high-frequency and geometry-independent texture.
By comparing Fig. 13(b)13(c) and 13(d) together, we can see that Detail Preserver effectively enhances the details of geometries. This insight is evident by this phenomenon: the frequency domain energy ratio between input and depth estimation w/ Detail Preserver, is closer to the frequency domain energy ratio between input and GT depth, compared with the frequency domain energy ratio between input and depth estimation w/o Detail Preserver.
Appendix G The effect of different time-steps $t$ in one-step diffusion
In Sec. 4.2 of our main paper, we reduce the number of training time-steps of diffusion formulation to only one, and fixing the only time-step $t$ to $T$ following the diffusion formulation. In this section, we evaluate the effect of different time-steps $t$ in one-step diffusion, rather than exclusively fixing $t=T$, to validate that the rule of basic diffusion formulations should better be followed. Violating it will lead to performance degradation. As shown in Tab. 8, we conduct experiments on Hypersim dataset (Roberts et al., 2021) and evaluated on NYUv2 dataset (Silberman et al., 2012), without employing the detail preserver or mixture dataset training. The results indicate that the model performs best when $t=T$ ($t=1000$). Changing $t$ leads to a slight degradation in performance.
Table 8: The effect of different time-steps $t$ in one-step diffusion. In this experiment, the models are trained on Hypersim dataset (Roberts et al., 2021) and evaluated on NYUv2 dataset (Silberman et al., 2012), without employing the detail preserver or mixture dataset training.
Time-step $t=1000$ $t=750$ $t=500$ $t=250$ $t=1$
AbsRel $\downarrow$ 5.587 5.631 5.727 5.663 5.737
$\delta 1$uparrow 96.272 96.165 96.087 96.141 96.080
@@ Appendix G The effect of different time-steps $t$ in one-step diffusion
Appendix H Qualitative Comparisons
@@
Appendix I Applications of Lotus
@@
Appendix J Future Work
@@
Some long invisible content to fill the page.
Another long invisible content to fill the page.